# Sampling 3D Molecular Conformers with Diffusion Transformers

### v1

@@ D.2 Input Featurization
D.3 Evaluation Metrics
During evaluation, we follow the same procedure as described in Refs. [65, 47, 46, 32]. The root-mean-square deviation (RMSD) metric measures the average distance between atoms of a generated conformer with respect to its reference, while taking into account all possible symmetries. For $L=2K$ let $\{\hat{C}_{l}\}_{l\in\{1,\dots,L\}}$ and $\{C_{k}\}_{k\in\{1,\dots,K\}}$ be the sets of generated conformers and reference conformers respectively. The Average Minimum RMSD (AMR) and Coverage (COV) metrics for both Recall (R) and Precision (P) are defined as follows, where $\delta>0$ is the coverage threshold:
@@ D.3 Evaluation Metrics
[…]
D.4 Chirality Correction
Given the four 3D coordinates around a chirality center denoted as $\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}}\in\mathbb{R}^{3}$ with $\mathbf{p_{i}}=(x_{i},y_{i},z_{i})$ for $i=1,2,3,4$, we can compute the volume $V$ of the tetrahedron as follows
$V(\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}})=\det\left(\begin{bmatrix}1&1&1&1\\ x_{1}&x_{2}&x_{3}&x_{4}\\ y_{1}&y_{2}&y_{3}&y_{4}\\ z_{1}&z_{2}&z_{3}&z_{4}\end{bmatrix}\right)=(\mathbf{p_{1}}-\mathbf{p_{4}})\cdot((\mathbf{p_{2}}-\mathbf{p_{4}})\times(\mathbf{p_{3}}-\mathbf{p_{4}}))$ (A52)
@@ D.4 Chirality Correction
Following GeomMol [65] and ET-Flow [32] we can then compare the orientation of the volume given by $\text{sign}(V)$ with the local chirality label produced by RDKit, which corresponds to a certain orientation as well (CW = +1 and CCW = -1) [69]. If the orientation of the volume differs from the RDKit label, we correct the chirality of the conformer by reflecting its positions against the $z$-axis.
Appendix E Additional Experimental Results
E.1 Scaling Results for PE(3)
Figure A4: (A) Number of gradient steps vs. wall time for PE(3) and aPE on the GEOM-DRUGS dataset measured on Nvidia H100 GPU with batch size 128. (B) Coverage precision and recall as function of threshold $\delta$ and (C) mean AMR metrics for base (“B”) and large (“L”) model of DiTMC+PE(3).
We investigate the scaling behavior of our equivariant models in Fig. A4. Tab. A10 shows a comparison of our base and large models for GEOM drugs. Similar to its non-equivariant counterpart, our scaled PE(3)-L model yields improved results for precision metrics, while we observe only minor improvements for recall. Tab. A11, Tab. A9 and Tab. A13 further demonstrate improvements for ensemble property prediction and out-of-distribution generalization using the larger model.
E.2 Conditioning ablation
To evaluate the effectiveness of various graph conditioning strategies in DiTMC , we compare the performance of three conditioning methods – “node only”, “node & bond-pair”, and “node & all-pair” – against a baseline model without any conditioning. As reported in Tab. A8, all conditioning variants significantly reduce the Average Minimum RMSD (AMR) for both recall (AMR-R) and precision (AMR-P) across the aPE and PE(3) settings, compared to the unconditioned baseline.
Notably, the “node & all-pair” strategy achieves the best overall performance, with the lowest AMR values. These results highlight the strength of the all-pair conditioning strategy, which leverages graph geodesics to incorporate information from all atom pairs, rather than restricting conditioning to directly connected nodes or bonded pairs. This comprehensive approach captures more global structural information, thereby improving both precision and recall. See Appendix K for a more in-depth analysis.
Table A8: Ablation of conditioning strategies in DiTMC (in brackets) with absolute PE and PE(3) on GEOM-QM9. -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds with standard deviation reported below.
Mean Median Mean Median Mean Median Mean Median
DiTMC+PE(3)-B (no conditioning) 71.5 $\pm$0.5 99.0 $\pm$1.4 0.358 $\pm$0.003 0.266 $\pm$0.005 40.4 $\pm$0.2 42.2 $\pm$0.4 0.683 $\pm$0.003 0.653 $\pm$0.005
DiTMC+PE(3)-B (node only) 95.5 $\pm$0.4 100.0 $\pm$0.0 0.074 $\pm$0.003 0.026 $\pm$0.001 91.1 $\pm$0.4 100.0 $\pm$0.0 0.114 $\pm$0.003 0.041 $\pm$0.001
DiTMC+PE(3)-B (node & bond-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.001 0.022 $\pm$0.001 93.6 $\pm$0.2 100.0 $\pm$0.0 0.091 $\pm$0.003 0.035 $\pm$0.001
DiTMC+PE(3)-B (node & all-pair) 95.7 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.002 0.021 $\pm$0.001 93.4 $\pm$0.2 100.0 $\pm$0.0 0.089 $\pm$0.002 0.032 $\pm$0.001
E.3 Ensemble Properties
@@
While RMSD-based metrics offer a geometric perspective on the quality of the generated conformers, they do not assess the chemical fidelity. To address this, we evaluate the median averaged errors of different ensemble properties between the generated and ground truth conformers. We adopt the property prediction task setup from MCF [47] and ET-Flow [32], where we draw a subset of 100 randomly sampled molecules from the test set of GEOM-DRUGS and generate $\min(2K,32)$ conformers for a molecule with $K$ ground truth conformers. Afterwards we relax the conformers using GFN2-xTB [70] and compare the Boltzmann-weighted properties of the generated and ground truth ensembles. More specifically, we employ xTB [70] to calculate the energy $E$, the dipole moment $\mu$, the HOMO-LUMO gap $\Delta\epsilon$ and the minimum energy $E_{\text{min}}$. We repeat this procedure for three subsets each sampled with a different random seed and report the averaged median absolute error and standard deviation of the different ensemble properties in Table A11.
@@
Table A9: Results on GEOM-QM9 for different generative models (number of parameters in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
@@
Table A10: Results on GEOM-DRUGS for different generative models (number of parameters in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
@@
*DiTMC+PE(3)-L (28.2M) 80.8 $\pm$0.3 85.6 $\pm$0.1 0.415 $\pm$0.003 0.376 $\pm$0.001 76.4 $\pm$0.2 82.6 $\pm$0.3 0.491 $\pm$0.002 0.414 $\pm$0.004
Table A11: Median absolute error of ensemble properties between generated and reference conformers. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Results for MCF, ET-Flow, and ours are averaged over three random seeds.
@@
OMEGA 0.68 0.66 0.68 0.69
GeoDiff 0.31 0.35 0.89 0.39
GeoMol 0.42 0.34 0.59 0.40
Torsional Diff. 0.22 0.35 0.54 0.13
MCF 0.68 $\pm$0.06 0.28 $\pm$0.05 0.63 $\pm$0.05 0.04 $\pm$0.00
ET-Flow 0.18 $\pm$0.01 0.18 $\pm$0.01 0.35 $\pm$0.06 0.02 $\pm$0.00
*DiTMC+aPE-B 0.17 $\pm$0.00 0.16 $\pm$0.01 0.27 $\pm$0.01 0.01 $\pm$0.00
*DiTMC+aPE-L 0.16 $\pm$0.02 0.14 $\pm$0.03 0.27 $\pm$0.01 0.01 $\pm$0.00
*DiTMC+PE(3)-B 0.18 $\pm$0.01 0.18 $\pm$0.01 0.27 $\pm$0.03 0.02 $\pm$0.00
*DiTMC+PE(3)-L 0.17 $\pm$0.01 0.14 $\pm$0.01 0.25 $\pm$0.01 0.01 $\pm$0.00
E.4 Generalization Results on GEOM-QM9
Additionally, we want to assess how well our proposed model architecture generalizes to unseen molecules. Here we evaluate our models trained on GEOM-DRUGS on the GEOM-QM9 dataset. We report the generalization performance in Table A12.
@@ E.4 Generalization Results on GEOM-QM9
Table A12: Out-of-distribution generalization results on GEOM-QM9 for models trained on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
@@
ET-Flow 86.7 100.0 0.218 0.160 68.7 75.3 0.369 0.317
*DiTMC+aPE-B 84.3 $\pm$0.5 100.0 $\pm$0.0 0.209 $\pm$0.004 0.134 $\pm$0.003 69.4 $\pm$0.7 84.4 $\pm$3.2 0.329 $\pm$0.004 0.250 $\pm$0.008
*DiTMC+aPE-L 80.6 $\pm$0.5 100.0 $\pm$0.0 0.199 $\pm$0.004 0.115 $\pm$0.004 68.83 $\pm$0.2 86.2 $\pm$2.0 0.329 $\pm$0.002 0.224 $\pm$0.007
*DiTMC+PE(3)-B 84.8 $\pm$0.3 100.0 $\pm$0.0 0.205 $\pm$0.003 0.130 $\pm$0.005 69.1 $\pm$0.1 81.41 $\pm$2.9 0.329 $\pm$0.004 0.254 $\pm$0.015
*DiTMC+PE(3)-L 85.1 $\pm$0.2 100.0 $\pm$0.0 0.195 $\pm$0.004 0.115 $\pm$0.003 69.5 $\pm$0.6 82.4 $\pm$2.3 0.319 $\pm$0.05 0.245 $\pm$0.009
E.5 Generalization Results on GEOM-XL
We also study how well our proposed model architecture generalizes to unseen molecules with a large number of atoms. For this we adopt the GEOM-XL dataset containing a total of 102 molecules with more than 100 atoms presented in [46]. We report the generalization performance in Table A13.
Table A13: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
Method AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ # mols
Mean Median Mean Median
GeoDiff 2.92 2.62 3.35 3.15 -
GeoMol 2.47 2.39 3.30 3.14 -
Tor. Diff. 2.05 1.86 2.94 2.78 -
MCF - S 2.22 1.97 3.17 2.81 102
MCF - B 2.01 1.70 3.03 2.64 102
MCF - L 1.97 1.60 2.94 2.43 102
ET-Flow 2.31 1.93 3.31 2.84 102
*DiTMC+aPE-B 1.96 $\pm$0.00 1.60 $\pm$0.03 2.90 $\pm$0.00 2.48 $\pm$0.03 102
*DiTMC+aPE-L 1.88 $\pm$0.01 1.51 $\pm$0.02 2.81 $\pm$0.00 2.30 $\pm$0.02 102
*DiTMC+PE(3)-B 1.98 $\pm$0.01 1.67 $\pm$0.02 3.03 $\pm$0.00 2.60 $\pm$0.01 102
*DiTMC+PE(3)-L 1.85 $\pm$0.02 1.58 $\pm$0.03 2.93 $\pm$0.00 2.53 $\pm$0.03 102
Tor. Diff. 1.93 1.86 2.84 2.71 77
MCF - S 2.02 1.87 2.9 2.69 77
MCF - B 1.71 1.61 2.69 2.44 77
MCF - L 1.64 1.51 2.57 2.26 77
ET-Flow 2.00 1.80 2.96 2.63 75
*DiTMC+aPE-B 1.68 $\pm$0.00 1.47 $\pm$0.02 2.59 $\pm$0.00 2.24 $\pm$0.01 77
*DiTMC+aPE-L 1.56 $\pm$0.01 1.28 $\pm$0.01 2.47 $\pm$0.00 2.14 $\pm$0.01 77
*DiTMC+PE(3)-B 1.73 $\pm$0.01 1.55 $\pm$0.01 2.71 $\pm$0.00 2.35 $\pm$0.01 77
*DiTMC+PE(3)-L 1.57 $\pm$0.01 1.46 $\pm$0.01 2.60 $\pm$0.00 2.27 $\pm$0.02 77
Appendix F Additional Ablations
F.1 Index Positional Encoding (iPE)

### v2

@@ G.2 Input Featurization
G.3 Evaluation Metrics
During evaluation, we follow the same procedure as described in Refs. [ganea2021geomol, wang2023swallowing, jing2022torsional, hassan2024etflow]. The root-mean-square deviation (RMSD) metric measures the average distance between atoms of a generated conformer with respect to its reference, while taking into account all possible symmetries. For $L=2K$ let $\{\hat{C}_{l}\}_{l\in\{1,\dots,L\}}$ and $\{C_{k}\}_{k\in\{1,\dots,K\}}$ be the sets of generated conformers and reference conformers respectively. The average minimum RMSD (AMR) and coverage (COV) metrics for both recall (R) and precision (P) are defined as follows, where $\delta>0$ is the coverage threshold:
$\text{COV-R}(C,\hat{C},\delta)$ $:=\frac{1}{K}\left|\left\{k\in\{1,\dots,K\}\mid\exists_{l\in\{1,\dots,L\}}\text{RMSD}(\hat{C}_{l},C_{k})<\delta\right\}\right|$ (A49)
$\text{COV-P}(C,\hat{C},\delta)$ $:=\frac{1}{L}\left|\left\{l\in\{1,\dots,L\}\mid\exists_{k\in\{1,\dots,K\}}\text{RMSD}(\hat{C}_{l},C_{k})<\delta\right\}\right|$ (A50)
@@ G.3 Evaluation Metrics
G.4 Chirality Correction
Given the four 3D coordinates around a chirality center denoted as $\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}}\in\mathbb{R}^{3}$ with $\mathbf{p_{i}}=(x_{i},y_{i},z_{i})$ for $i=1,2,3,4$, we can compute the oriented volume $OV$ of the tetrahedron via
@@ G.4 Chirality Correction
$=\frac{1}{6}\cdot(\mathbf{p_{1}}-\mathbf{p_{4}})\cdot((\mathbf{p_{2}}-\mathbf{p_{4}})\times(\mathbf{p_{3}}-\mathbf{p_{4}}))\,.$
@@
Following GeomMol [ganea2021geomol] and ET-Flow [hassan2024etflow], we can then compare the orientation of the volume given by $\text{sign}(OV)$ with the local chirality label produced by RDKit, which corresponds to a certain orientation as well (CW = +1 and CCW = -1) [pattanaik2020message]. If the orientation of the volume differs from the RDKit label, we correct the chirality of the conformer by reflecting its positions against the $z$-axis.
G.5 Pareto Front
For all models, we generate conformers using 5, 10, 20 and 50 sampling steps on a single A100 GPU with a batch size of 128, following Refs. [hassan2024etflow, wang2023swallowing]. The wall-clock time per generated sample is obtained by measuring the average time per batch and dividing by the batch size. As done in the original paper, we adopt DDIM sampling for MCF-S, MCF-B and MCF-L.
G.6 Ensemble Properties
@@
We adopt the property prediction task setup from MCF [wang2023swallowing] and ET-Flow [hassan2024etflow], where we draw a subset of 100 randomly sampled molecules from the test set of GEOM-DRUGS and generate $\min(2K,32)$ conformers for a molecule with $K$ ground truth conformers. Afterwards we relax the conformers using GFN2-xTB [bannwarth2019xtb] and compare the Boltzmann-weighted properties of the generated and ground truth ensembles. More specifically, we employ xTB [bannwarth2019xtb] to calculate the energy $E$, the dipole moment $\mu$, the HOMO–LUMO gap $\Delta\epsilon$ and the minimum energy $E_{\text{min}}$. We repeat this procedure for three subsets each sampled with a different random seed and report the averaged median absolute error and standard deviation of the different ensemble properties.
Appendix H Additional Experimental Results
H.1 Results on GEOM-QM9
We report the results on GEOM-QM9 including standard deviations for all DiTMC models in Tab. A9.
@@ H.1 Results on GEOM-QM9
Table A9: Results on GEOM-QM9 for different generative models (number of parameters in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
@@
H.2 Results on GEOM-DRUGS
We report the results on GEOM-DRUGS including standard deviations for all DiTMC models in Tab. A10. The median absolute error of ensemble properties on GEOM-DRUGS is shown in Tab. A11.
Additional results for the coverage vs. RMSD threshold analysis, including results for DiTMC+rPE and DiTMC+PE(3), are provided in Fig. A8, Fig. A9, and Fig. A10.
Additional results for the Pareto front analysis, including results for DiTMC+rPE and DiTMC+PE(3), are provided in Fig. A11, Fig. A12, and Fig. A13.
Table A10: Results on GEOM-DRUGS for different generative models (number of parameters in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
@@ H.2 Results on GEOM-DRUGS
*DiTMC+aPE-L (28.2M) 79.2 $\pm$0.1 84.4 $\pm$0.2 0.432 $\pm$0.003 0.386 $\pm$0.003 77.8 $\pm$0.1 85.7 $\pm$0.5 0.470 $\pm$0.001 0.387 $\pm$0.003
*DiTMC+rPE-L (28.3M) 78.7 $\pm$0.1 84.1 $\pm$0.4 0.438 $\pm$0.002 0.388 $\pm$0.005 78.1 $\pm$0.1 86.4 $\pm$0.3 0.466 $\pm$0.001 0.381 $\pm$0.003
*DiTMC+PE(3)-L (31.1M) 80.8 $\pm$0.3 85.6 $\pm$0.1 0.415 $\pm$0.003 0.376 $\pm$0.001 76.4 $\pm$0.2 82.6 $\pm$0.3 0.491 $\pm$0.002 0.414 $\pm$0.004
Table A11: Median absolute error of ensemble properties between generated and reference conformers. Best results in bold, second best underlined; our models are marked with an asterisk “*”. Results for MCF, ET-Flow, and ours are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
@@
GeoDiff (1.6M) 0.31 0.35 0.89 0.39
GeoMol (0.3M) 0.42 0.34 0.59 0.40
Torsional Diff. (1.6M) 0.22 0.35 0.54 0.13
MCF-L (242M) 0.68 $\pm$0.06 0.28 $\pm$0.05 0.63 $\pm$0.05 0.04 $\pm$0.00
ET-Flow (8.3M) 0.18 $\pm$0.01 0.18 $\pm$0.01 0.35 $\pm$0.06 0.02 $\pm$0.00
*DiTMC+aPE-B (9.5M) 0.17 $\pm$0.00 0.16 $\pm$0.01 0.27 $\pm$0.01 0.01 $\pm$0.00
*DiTMC+aPE-L (28.2M) 0.16 $\pm$0.02 0.14 $\pm$0.03 0.27 $\pm$0.01 0.01 $\pm$0.00
*DiTMC+rPE-B (9.6M) 0.16 $\pm$0.03 0.16 $\pm$0.03 0.29 $\pm$0.02 0.02 $\pm$0.00
*DiTMC+rPE-L (28.3M) 0.16 $\pm$0.01 0.15 $\pm$0.02 0.28 $\pm$0.06 0.01 $\pm$0.00
*DiTMC+PE(3)-B (8.6M) 0.18 $\pm$0.01 0.18 $\pm$0.01 0.27 $\pm$0.03 0.02 $\pm$0.00
*DiTMC+PE(3)-L (31.1M) 0.17 $\pm$0.01 0.14 $\pm$0.01 0.25 $\pm$0.01 0.01 $\pm$0.00
H.3 Results on GEOM-XL
We report the results on GEOM-XL including standard deviations for all DiTMC models in Tab. A12.
@@ H.3 Results on GEOM-XL
Table A12: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
Method 75 / 77 mols 102 mols
AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
Tor. Diff. (1.6M) 1.93 1.86 2.84 2.71 2.05 1.86 2.94 2.78
MCF-S (13M) 2.02 1.87 2.90 2.69 2.22 1.97 3.17 2.81
MCF-B (64M) 1.71 1.61 2.69 2.44 2.01 1.70 3.03 2.64
MCF-L (242M) 1.64 1.51 2.57 2.26 1.97 1.60 2.94 2.43
ET-Flow (8.3M) 2.00 1.80 2.96 2.63 2.31 1.93 3.31 2.84
*DiTMC+aPE-B (9.5M) 1.68 $\pm$0.00 1.47 $\pm$0.02 2.59 $\pm$0.00 2.24 $\pm$0.01 1.96 $\pm$0.00 1.60 $\pm$0.03 2.90 $\pm$0.00 2.48 $\pm$0.03
*DiTMC+aPE-L (28.2M) 1.56 $\pm$0.01 1.28 $\pm$0.01 2.47 $\pm$0.00 2.14 $\pm$0.01 1.88 $\pm$0.01 1.51 $\pm$0.02 2.81 $\pm$0.00 2.30 $\pm$0.02
*DiTMC+rPE-B (9.6M) 1.69 $\pm$0.01 1.41 $\pm$0.03 2.52 $\pm$0.00 2.11 $\pm$0.00 1.97 $\pm$0.01 1.61 $\pm$0.01 2.86 $\pm$0.00 2.33 $\pm$0.01
*DiTMC+rPE-L (28.3M) 1.66 $\pm$0.03 1.37 $\pm$0.01 2.47 $\pm$0.00 2.18 $\pm$0.02 1.96 $\pm$0.02 1.61 $\pm$0.02 2.82 $\pm$0.00 2.42 $\pm$0.02
*DiTMC+PE(3)-B (8.6M) 1.73 $\pm$0.01 1.55 $\pm$0.01 2.71 $\pm$0.00 2.35 $\pm$0.01 1.98 $\pm$0.01 1.67 $\pm$0.02 3.03 $\pm$0.00 2.60 $\pm$0.01
*DiTMC+PE(3)-L (31.1M) 1.57 $\pm$0.01 1.46 $\pm$0.01 2.60 $\pm$0.00 2.27 $\pm$0.02 1.85 $\pm$0.02 1.58 $\pm$0.03 2.93 $\pm$0.00 2.53 $\pm$0.03
Appendix I Additional Ablations
I.1 Gaussian vs. Harmonic Prior
As shown in Tab. A13, using the harmonic prior improves all metrics slightly for our models on GEOM-QM9. Using the harmonic prior however doesn’t seem to be a crucial ingredient for the success of our method, as differences between Gaussian and Harmonic prior appear diminishing. As the results in Tab. A13 verify, our method can also be used with a simple Gaussian prior effectively. For larger molecular graphs the expensive eigendecomposition of the graph Laplacian required for the Harmonic prior could therefore be avoided, which helps scaling our approach more easily.
Table A13: Ablation of PE strategies and Gaussian (G) and Harmonic (H) prior on GEOM-QM9. We report mean coverage (COV) at a threshold of $0.5\mbox{\AA }$, and mean average minimum RMSD (AMR) for Recall -R and Precision -P. Best results in bold. All results are averaged over three random seeds.
G H G H G H G H
DiTMC+aPE-B 96.2 96.1 0.074 0.073 95.2 95.4 0.087 0.085
DiTMC+rPE-B 96.0 96.3 0.073 0.070 95.2 95.7 0.084 0.080
DiTMC+PE(3)-B 95.7 95.7 0.069 0.068 93.5 93.4 0.090 0.089
I.2 Conditioning Strategies on GEOM-QM9
To evaluate the effectiveness of various graph conditioning strategies in DiTMC , we compare the performance of different conditioning methods against a baseline model without any conditioning. In addition to conditioning strategies discussed in Sec. 4.1, we note that our conditioning GNN also produces edge-level representations, which can be used to define pair-wise graph conditioning tokens as
$\textbf{bond-pair:}\quad\boldsymbol{c}_{ij}^{\mathcal{G}}=\begin{cases}\text{GNN}_{\text{edge}}(\mathcal{V},\mathcal{E})&\forall(i,j)\in\mathcal{E}\\ \boldsymbol{\bar{c}}^{\mathcal{G}}&\forall(i,j)\not\in\mathcal{E}\,.\end{cases}$ (A54)
These tokens only capture interactions between bonded atoms, i.e., when $(i,j)\in\mathcal{E}$. Conditioning tokens for non-bonded pairs are set to a learnable vector $\boldsymbol{\bar{c}}^{\mathcal{G}}$. Self-attention still operates on all atom pairs $(i,j)$, even if they are not connected by a chemical bond.
Specifically, we ablate the following conditioning strategies:
  • node only conditioning using only atom-wise graph conditioning tokens $\boldsymbol{c}_{i}^{\mathcal{G}}$ (Eq. 6).
  • node & bond-pair conditioning using both atom-wise graph conditioning tokens $\boldsymbol{c}_{i}^{\mathcal{G}}$ (Eq. 6) and pair-wise graph conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}$ derived from edge-level representations of the conditioning GNN as discussed above (Eq. A54).
  • node & all-pair conditioning using both atom-wise graph conditioning tokens $\boldsymbol{c}_{i}^{\mathcal{G}}$ (Eq. 6) and pair-wise graph conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}$ based on geodesic graph distances (Eq. 7).
As reported in Tab. A14, all our proposed conditioning strategies significantly reduce the average minimum RMSD (AMR) for both recall (AMR-R) and precision (AMR-P) using DiTMC+aPE-B, compared to the unconditioned baseline.
Notably, the “node & all-pair” strategy achieves the best overall performance, with the lowest AMR values. These results highlight the strength of the all-pair conditioning strategy, which leverages graph geodesics to incorporate information from all atom pairs, rather than restricting conditioning to directly connected nodes or bonded pairs. This comprehensive approach captures more global structural information, thereby improving both precision and recall. See Appendix L for a more in-depth analysis.
Table A14: Ablation of conditioning strategies using DiTMC+aPE-B on GEOM-QM9. -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds with standard deviation reported below.
@@ I.2 Conditioning Strategies on GEOM-QM9
I.3 Index Positional Encoding (iPE)